OpenStudy (anonymous):

Let be a random variable with possible values of 1,2,3,4,5 and pdf P(x)x^2/55. Find the cumulative distribution function for X.

OpenStudy (anonymous):

A probability and statistic problem

OpenStudy (anonymous):

If \(p(x)\) denotes the pdf, \(P(x)\) denotes the cdf, and \(\mathbb{P}(x=k)\) the probability that \(x\) takes value \(k\), then \[\large P(x)=\mathbb{P}(X\le k)=\sum_{x=-\infty}^kp(x)\] You'll have 7 cases to consider. For all \(x<1\), you have a cumulative probability because (presumably) the pdf is defined to be \(\dfrac{x^2}{55}\) for \(x=1,...,5\), and 0 elsewhere: \[\large \begin{align*}P(x)&=\begin{cases}\displaystyle\sum_{x=-\infty}^00&\text{for }x<1\end{cases}\\\\ &= \begin{cases} 0&\text{for }x<1 \end{cases}\end{align*}\] For \(x=1\), you have a cumulative probability of \[\large \begin{align*}P(x)&=\begin{cases}\displaystyle\sum_{x=-\infty}^00&\text{for }x<1\\\\ \displaystyle\sum_{x=-\infty}^1\frac{x^2}{55}&\text{for }x=1\end{cases}\\\\ &= \begin{cases} 0&\text{for }x<1\\\\ \dfrac{1}{55}&\text{for }x=1 \end{cases}\end{align*}\] For \(x=2\), you have \[\large \begin{align*}P(x)&=\begin{cases}\displaystyle\sum_{x=-\infty}^00&\text{for }x<1\\\\ \displaystyle\sum_{x=-\infty}^1\frac{x^2}{55}&\text{for }x=1\\\\ \displaystyle\sum_{x=-\infty}^2\frac{x^2}{55}&\text{for }x=2\end{cases}\\\\ &= \begin{cases} 0&\text{for }x<1\\\\ \dfrac{1}{55}&\text{for }x=1\\\\ \dfrac{1}{55}+\dfrac{4}{55}&\text{for }x=2 \end{cases}\end{align*}\] and so on, up to the case where \(x>5\).

OpenStudy (anonymous):

thank you much i was so loss about that problem lol